# Coupled resonators based on high permittivity dielectrics for microwave sensors

### v1

@@ III Coupled STO and elliptical cavity
In practice, eigenmode information is obtained through S-parameter measurements. Figure 7 illustrates the S-parameters for an excitation scheme using two antennas (as shown in Fig. 4), with a puck offset of 50 mm. Two key parameters in the measurement are the frequency sensitivity of the modes to variations in $\varepsilon_{r}$ and the $Q$-factor of the modes. The quality factor can be calculated using a 3 dB method and is also related to phase variations in the $S_{1,2}$ curve; larger phase changes (i.e., a steeper phase derivative) correspond to a higher $Q$-factor. Using the phase derivative offers the benefit of not only analyzing the peaks but also assessing the sensitivity of the notch. This notch, located between two resonance peaks in the S-parameters, corresponds to a point of minimal signal transmission.
@@
Figure 8 presents the frequencies of the two peaks in the $S_{1,2}$ curve, along with the notch between them. The resonance behavior as a function of $\varepsilon_{r}$ resembles eigenmode analysis, with level repulsion influenced by the $r_{\mathrm{offset}}$. The frequency of the notch shifts similarly to that of the single STO puck calculated from the semi-analytical formula
$f\mathrm{[GHz]}\approx\frac{34}{a\mathrm{[mm]}\sqrt{\varepsilon_{\mathrm{r}}}}\left(\frac{a}{d}+3.45\right)$ (1)
where $a$ and $d$ are the radius and height of the puck, respectively [38]. The frequency sensitivity of modes 1 and 2 is generally lower than that of the notch as shown in Fig. 8(b). Similar to the eigenmode analysis, for $\varepsilon_{r}$ below 230, mode 1 is the elliptical cavity-dominant mode, whereas above 230, it becomes the second mode. For the elliptical cavity-dominant mode, frequency sensitivity peaks at $\varepsilon_{r}=230$, although this is accompanied by a decrease in its phase derivative amplitude that is related to $Q$-factor value. However, the product of frequency sensitivity and the phase derivative for the two modes is equal to that of the notch for a PEC cavity (Fig. 8(c)), whereas for a copper cavity (Fig. 8(d)), this product for the notch exceeds that of the two modes. This coupling allows for a trade-off between higher frequency sensitivity and a higher $Q$-factor, depending on the $\varepsilon_{r}$ value.
(a) Measurement setup of the coupled cavity
(b) Measurement results
(c) Measurement results over a wide frequency range
Figure 9: The measurement setup and results of the coupled elliptical cavity with the STO puck. The STO puck is fixed to the flange via a long ceramic rod at $r_{\mathrm{offset}}$=19 mm. Two long antennas were used for the $S_{2,1}$ measurement: one at the flange center and the other at $r_{\mathrm{offset}}$=-19 mm. Note that the elliptical cavity used for the measurements had a slightly different shape than the one used for the simulations, causing a frequency shift in the resonance modes of the cavity. An $\varepsilon_{\mathrm{r}}$ of 300 is assumed in the simulation.
@@
To assess the coupling between the two modes in practice, a test case was established by placing the STO puck inside an elliptical cavity. Excitation is provided through two long rods positioned at the end flanges, and the STO puck is secured at the center of the cavity with a radial offset of 19 mm, mounted on a ceramic rod (see Fig. 9). The resonance of the STO puck, observed at 1.16 GHz in Fig. 9(b), aligns with the predictions from the simulations. This correspondence suggests that the analyses conducted in this section regarding the coupling between the two modes can be considered valid. In addition to the fundamental mode of the STO puck and the elliptical cavity, many other higher-order modes can also be excited in the resonators (as shown in Fig. 9(c)). Depending on the application, one or more of these modes with different frequencies, field patterns, and frequency sensitivities could also be selected for mode coupling.
@@ IV Discussion
Coupled dielectric resonators have already been proposed for the realization of sensitive particle and radiation bolometers at frequency around 10 GHz. In the bolometer detector described in Ref. [21, 22, 23], two resonators, realized using Sapphire and STO pucks, are positioned one on the top of the other. Their coupling by fringing fields turns out to be very weak, thus they can be described by using a perturbative theory. Fixed frequency bias of the system then results in a potentially large change in the output amplitude signal at the working frequency. The energy resolution that could be achieved is $\delta E\sim 10^{-23}$ J/Hz at 1 K, $10^{-20}$ J/Hz at 10 K and $2.5\times 10^{-19}$  J/Hz at 30 K. Unlike other high-resolution detectors, the bolometer operates at a relatively high cryogenic temperature in the range between 7 and 30 K. In addition to the detection of single particles, a rapid thermal time constant and a high $df/dT$ response have been demonstrated [21, 22, 23].
In our case, the sapphire resonator is replaced with a superconducting cavity and, as a paradigmatic example, we have used an elliptical TESLA-shaped cavity available at CERN. When the STO resonator is positioned inside the cavity, the coupling between the two can be easily controlled by varying the puck’s offset and frequency (Sect. III). In particular, the latter can be defined by properly adjusting the physical dimensions of the puck and by taking into account the STO permittivity at the selected working temperature. In this regard, it is important to stress that the results reported in Fig. 5, 7 and 8, which are shown as a function of the STO permittivity and by assuming the geometric parameters of the STO puck described in Sec. II, are valid also for higher values of the STO permittivity such as those found at low temperature.
Given the remarkable change of the STO permittivity as a function of the temperature, we expect that resonators having typical diameter in the range of 1 mm are needed to achieve a significant coupling with the TESLA-shaped cavity at liquid helium temperatures, or below. On the one hand, the heat capacity of the puck, which is directly related to its volume, plays a significant role in the system’s effectiveness as a detector. Reducing the puck’s size increases sensitivity by amplifying the temperature rise per unit of absorbed energy, thereby enhancing the system’s detection capabilities. On the other hand, the coupling may be significant reduced with a small STO puck. The exploitation of higher microwave modes of the dielectric resonator may constitute an useful resource in this context. Alternative dielectrics have low temperature permittivity values nearer to those considered in the simulations; in particular we have considered rutile resonators, which at 0.8 K show $Q$-factor of $5.6\times 10^{5}$ for the mode at 2.6 GHz [40]. However the variation of the permittivity of rutile as a function of temperature is less marked [41], thus coupled resonator systems based on rutile resonators are expected to be less sensitive. Different kinds of superconducting cavities can also be considered to operate at different frequencies.
@@
From the simulations reported in Fig. 5 it results that the two cavities can be tuned either in a strong or weak coupling regime. As expected, at resonance the quality factor of the superconducting cavity progressively decreases as the coupling increases. However, we note that eigenmodes having $Q$-factor in the $10^{5}-10^{6}$ range can still be obtained by adjusting $r_{\mathrm{offset}}$ and the detuning of the natural frequencies of superconducting cavity and STO resonator. These values are corroborated by transmission spectra measured at low temperature (Fig. 2), which show that the $Q$-factor of the STO resonator is in the $10^{4}$ range. This value can potentially be further improved by using opportunely prepared STO pucks. Electromagnetic simulations also show that the frequency and phase shifts of the superconducting cavity are amplified as an effect of the coupling with the STO resonator (Fig. 7). In this regard, the phase derivative of $S_{1,2}$ show a particularly steep response.
@@
We also mention that preliminary tests were carried out to test the stability of the STO mode frequency. In the case of measurements running for hours, the observed drift of the resonance frequency, due to pexternal factors and without any feedback control, was in the range of 50 Hz both for the STO puck measured at 0.1 K and for superconducting cavities measured at 1.7 K. These values are strongly reduced, of about one order of magnitude, for timescales of few seconds or minutes. By improving the experimental set up, and introducing a reference phase-locked loop (PLL) frequency stabilization, it is expected a significant reduction of the frequency drift, thus allowing to exploit phase measurements and further improve the system sensitivity. Additionally, the frequency tuning of the detection cavity can be ensured by the ferroelectricity of the STO and can be obtained by applying a DC electric field to the dielectric resonator. The expected tuning range is in the order of few percent.
@@ V Conclusions
To conclude, we have investigated the coupling between a high-$Q$ superconducting cavity and a STO resonator. This study is critical for ultimately optimizing the STO puck dimensions such that mode coupling occurs at the desired cold temperatures, where the STO’s frequency is most responsive to temperature changes. The dielectric constant can reach values higher than 30000 measured at temperature of 0.16 K and frequency of 0.1 GHz with a $Q$-factor of 10000, and can be very well coupled with a 1.3 GHz TESLA cavity, at least at room temperature. The simulations performed with the given STO dimensions indicate a frequency sensitivity of 2.8 MHz for the notch, corresponding to $\varepsilon_{\mathrm{r}}$ around 230, where the coupling is at its maximum. This sensitivity aligns with the pattern predicted by the semi-analytic formula for the STO puck alone. Consequently, this approach allows for precise calculation of the STO puck dimensions, providing the necessary frequency sensitivity for a specific project at the target frequency.
@@ VI ACKNOWLEDGMENTS
We thank Fritz Casper, Sergio Calatroni, Akira Miyazaki, Walter Venturini, Giovanni Carugno and Marco Affronte for stimulating discussions. A.G. acknowledge financial support from PNRR MUR project ECS_00000033_ECOSISTER and A.C. from PNRR MUR project E63C22002190007 , PE23 , NQSTI.

### v2

@@ III Coupled STO and elliptical cavity
In practice, eigenmode information is obtained through S-parameter measurements. Figure 7 illustrates the S-parameters for an excitation scheme using two antennas (as shown in Fig. 4), with a puck offset of 50 mm. Two key parameters in the measurement are the frequency sensitivity of the modes to variations in $\varepsilon_{r}$ and the $Q$-factor of the modes. The quality factor can be calculated using a 3 dB method and is also related to phase variations in the $S_{1,2}$ curve; larger phase changes (i.e., a steeper phase derivative) correspond to a higher $Q$-factor. Using the phase derivative offers the benefit of not only analyzing the peaks but also assessing the sensitivity of the transmission zero. This notch, located between two resonance peaks in the S-parameters, corresponds to a point of minimal signal transmission.
@@
Figure 8 presents the frequencies of the two peaks in the $S_{1,2}$ curve, along with the notch between them. The resonance behavior as a function of $\varepsilon_{r}$ resembles eigenmode analysis, with level repulsion influenced by the $r_{\mathrm{offset}}$. The frequency of the notch shifts similarly to that of the single STO puck calculated from the semi-analytical formula Eq. 1. The frequency sensitivity of modes 1 and 2 is generally lower than that of the notch as shown in Fig. 8(b). Similar to the eigenmode analysis, for $\varepsilon_{r}$ below 230, mode 1 is the elliptical cavity-dominant mode, whereas above 230, it becomes the second mode. For the elliptical cavity-dominant mode, frequency sensitivity peaks at $\varepsilon_{r}=230$, although this is accompanied by a decrease in its phase derivative amplitude that is related to $Q$-factor value. However, the product of frequency sensitivity and the phase derivative for the two modes is equal to that of the notch for a PEC cavity (Fig. 8(c)), whereas for a copper cavity (Fig. 8(d)), this product for the notch exceeds that of the two modes. This coupling allows for a trade-off between higher frequency sensitivity and a higher $Q$-factor, depending on the $\varepsilon_{r}$ value.
Figure 9: The measurement setup (a) and measurement results of the coupled elliptical cavity with the STO puck over a narrow frequency range (b) and a wide frequency range (c). The STO puck is fixed to the flange via a long ceramic rod at $r_{\mathrm{offset}}$=19 mm. Two long antennas were used for the $S_{2,1}$ measurement: one at the flange center and the other at $r_{\mathrm{offset}}$=-19 mm. Note that the elliptical cavity used for the measurements had a slightly different shape than the one used for the simulations, causing a small change of the resonance modes of the cavity. An $\varepsilon_{\mathrm{r}}$ of 300 is assumed in the simulation.
@@
To assess the coupling between the two modes in practice, a test case was established by placing the STO puck inside an elliptical cavity. Excitation is provided through two long rods positioned at the end flanges, and the STO puck is secured at the center of the cavity with a radial offset of 19 mm, mounted on a ceramic rod (see Fig. 9). The resonance of the STO puck, observed at 1.16 GHz in Fig. 9(b), is consistent with the predictions of the simulations, corroborating the analysis reported in this section. In addition to the fundamental mode of the STO puck and the elliptical cavity, many other higher-order modes can also be excited in the resonators, which are partially reproduced by the simulated model (Fig. 9(c)). Depending on the application, one or more of these modes with different frequencies, field patterns, and frequency sensitivities could also be selected for mode coupling.
@@ IV Discussion
Unlike the power detection scheme, the fundamental limit in frequency measurements comes from the best achievable frequency stability of the measurement parts and thus it is not directly related to the Nyquist noise and the ambient temperature, although cryogenic cooling may significantly improve the detection limit by increasing the quality factors. By using superconducting cavities and sapphire, it is possible to achieve $Q$-factors in the $10^{9}$ range [44] and bandwidths of the order of 1 Hz at microwave frequencies [45, 46]. This, also considering that the resolution of the resonator linewidth is typically limited, by electronic noise, to about 1 ppm [44], explains the unprecedented sensitivity of the frequency measurement approach.
By conjugating the frequency sensitivity of the high-$Q$ cavity and the tunability introduced by coupling it with the STO resonator (Sec. III), we can envisage the possible use of the hybrid system in sensing applications. Although the specific characteristics of a particular sensor depends by its geometry and working conditions, we can anticipate the potentialities of the proposed hybrid system in this context. In particular, coupled dielectric resonators have already been proposed for the realization of sensitive particle and radiation bolometers at frequency around 10 GHz. In the bolometer detector described in Ref. [12, 47, 20, 21, 22], two resonators, realized using Sapphire and STO pucks, are positioned one on the top of the other and coupled by fringing fields. In the weak coupling regime, the maximum sensitivity to frequency occurs when the frequency of the two resonators exactly coincides and [47]
$\left.\frac{\mathrm{d}f(T)}{\mathrm{d}T}\right|_{\mathrm{max}}=\frac{A}{(W_{\mathrm{STO}})^{2}}\frac{\mathrm{d}f_{\mathrm{STO}}}{\mathrm{d}T},$ (2)
being $A$ the coefficient that accounts for the coupling between the resonators modes and $W_{\mathrm{STO}}$ the unperturbed linewidth of STO [12]. The sensitivity of the bolometer is thus related to the temperature dependence of the frequency of the STO resonator ($\mathrm{d}f_{\mathrm{STO}}/\mathrm{d}T$). From Eq. 1 we obtain
$\frac{\mathrm{d}f_{\mathrm{STO}}}{\mathrm{d}T}=-\frac{\alpha}{2}\varepsilon_{\mathrm{r}}^{-3/2}\frac{\mathrm{d}\varepsilon_{\mathrm{r}}}{\mathrm{d}T},$ (3)
being $\alpha$ related to the dimensions of the STO puck. Eq. 2 shows that the maximum sensitivity can be amplified by properly arranging the $A/(W_{\mathrm{STO}})^{2}$ factor.
We focus here on the possibility to exploit the coupled cavity-STO resonator system as a bolometer, that is by replacing the originally proposed sapphire resonator with an elliptical cavity. The main advantages of this scheme are the possibility to tune the coupling strength by varying the STO puck position within the cavity and, in particular, to achieve higher couplings than those due to fringing fields of dielectric resonators positioned at a short distance [12]. As shown by electromagnetic simulations reported in Sec. III, the frequency sensitivity is quantified as $\mathrm{d}f/\mathrm{d}\varepsilon_{\mathrm{r}}=2.8\leavevmode\nobreak\ \mathrm{MHz}$ for $\varepsilon_{\mathrm{r}}\approx 230$. How this number is translated in $\mathrm{d}f/\mathrm{d}T$ sensitivity depends by the dimensions of the resonator and by the specific working conditions, in particular by the temperature range.
The temperature-dependence of the STO resonator frequency is not trivial owing to the structural transition around 100 K, the ferroelectric phase transition at 51 K and the stabilization of the quantum paraelectric phase below 5 K [35, 37, 36]. Let’s first consider temperatures between 20 and 80 K, where the relative permittivity of STO can be approximated as $\varepsilon_{\mathrm{r}}\approx 10^{5}/T\approx 2000$. This range was discussed in the original proposal by Gallop et al. [12] and the frequency sensitivity quantified as $\mathrm{d}f(T)/\mathrm{d}T=Af_{0}/(2W^{2}_{STO}T)$. Considering the outcome of simulations reported in Sec. III, we expect that higher couplings enable the improvement of $\mathrm{d}f(T)/\mathrm{d}T$ if the sapphire resonator is replaced by the metal cavity. Although the $Q$-factor of the normal-conducting cavity is typically lower than that of sapphire resonator [48], simulations suggest that, to obtain the maximum sensitivity, the change of the notch frequency should be considered (Fig. 8(d)).
We shall consider also the temperature range in which the cavity has superconducting properties. In particular, a Nb coated TESLA cavity is superconducting below 9.2 K where the $Q$-factor of the fundamental mode rises above $10^{9}$ [40]. Ideally, when the superconducting cavity is cooled below the critical temperature, the STO puck should be warm up to $T\approx 25\,\mathrm{K}$ to display the maximum value of $\mathrm{d}f_{\mathrm{STO}}/\mathrm{d}T$ [22]. This requirement, however, introduces further experimental complexity. Fig. 2 shows that between 1 and 0.4 K the frequency of the STO resonator alone decreases monotonically with temperature; from the fit we derive $\mathrm{d}f_{\mathrm{STO}}/\mathrm{d}T=10\,\mathrm{kHz/K}$ at 0.8 K. This number is much smaller than what obtained at $T\approx 20-25\,\mathrm{K}$ [22], however it is expected to increase if small STO pucks are used (Eq. 3), as required, when the permittivity rises up to $\varepsilon_{r}\approx 30000$, to match the frequencies of the STO resonator and cavity. Additionally, the hybrid system can be tuned to increase the coupling strength and compensate, at least partially, lower $\mathrm{d}f_{\mathrm{STO}}/\mathrm{d}T$ values in this temperature range (Eq. 2). Given $\varepsilon_{\mathrm{r}}\approx 30000$, we expect that resonators having typical diameter in the range of 1 mm are needed to achieve a significant resonant coupling with the 1.3 GHz cavity. Considering that the heat capacity of the puck, which is directly related to the mass, plays a significant role in the system’s effectiveness as a bolometer [47], we expect that reducing the puck’s size increases sensitivity by amplifying the temperature rise per unit of absorbed energy, thereby enhancing the system’s detection capabilities.
@@
Given the very peculiar characteristic of STO and the dependence of the STO resonator frequency by external factors such as applied electric field or mechanical stress, we just mention that possible applications of the proposed system of coupled cavity and STO puck are not limited to the realization of bolometers. Simulations indicate that the two cavities can be tuned to reach the strong coupling regime (Fig. 5). As expected, at resonance the quality factor of the superconducting cavity progressively decreases as the coupling increases. However, we note that eigenmodes having $Q$-factor in the $10^{5}-10^{6}$ range can still be obtained by adjusting $r_{\mathrm{offset}}$ and the detuning of the natural frequencies of superconducting cavity and STO resonator. These values are corroborated by transmission spectra measured at low temperature (Fig. 2), which show that the $Q$-factor of the STO resonator is in the $10^{4}$ range. This value can potentially be further improved by using opportunely prepared STO pucks. Electromagnetic simulations also show that the change of the frequency of the superconducting cavity modes location, and of the corresponding phase, are amplified as an effect of the coupling with the STO resonator (Fig. 7). In this regard, the phase derivative of $S_{1,2}$ shows a particularly steep response.
@@
We also mention that preliminary tests were carried out to test the stability of the STO mode frequency. In the case of measurements running for hours, the observed drift of the resonance frequency, due to external factors and without any feedback control, was in the range of 50 Hz both for the STO puck measured at 0.1 K and for superconducting cavities measured at 1.7 K. These values are strongly reduced, of about one order of magnitude, for timescales of few seconds or minutes. By improving the experimental set up, and introducing a reference phase-locked loop (PLL) frequency stabilization, it is expected a significant reduction of the frequency drift, thus allowing to exploit phase measurements and further improve the system sensitivity. Additionally, the frequency tuning of the detection cavity can be ensured by the ferroelectricity of the STO and can be obtained by applying a DC electric field to the dielectric resonator. The expected tuning range is in the order of few percent.
@@ V Conclusions
To conclude, we have studied the hybrid system composed by a high-$Q$ TESLA-shaped elliptical cavity and STO resonator, and investigated the effect of parameters, such as STO permittivity, puck dimensions and position within the cavity, that govern the coupling between the electromagnetic modes. Finite-element simulations show that the hybrid system offers great versatility to tune the coupling strength and achieve the strong coupling regime. These results are supported by test measurements carried out at room temperature using a copper cavity and a STO puck, and by the low temperature characterization of the STO resonator, which shows resonant frequency of 0.1 GHz and $Q$-factor of 10000 at 0.16 K. The hybrid system show potential for the realization of microwave sensors in which the sensitivity of the STO puck to selected physical quantities is exploited as the active element, while frequency or phase measurements on the high-$Q$ cavity are used to efficiently detect such changes. In particular, the application of the hybrid system in bolometers exploiting the temperature dependence of the STO permittivity has been discussed. Our results are useful to design dedicated experiments at low temperature allowing the direct test of sensitivity and tunability of the proposed hybrid system in high precision frequency-domain measurements.
@@ VI ACKNOWLEDGMENTS
We thank Fritz Casper, Sergio Calatroni, Akira Miyazaki, Walter Venturini, Alick Mcpherson, Giovanni Carugno and Marco Affronte for stimulating discussions. A.C. acknowledges financial support from PNRR MUR project E63C22002190007, PE23, NQSTI. A.G. acknowledges financial support from PNRR MUR project ECS_00000033_ECOSISTER. S.P. acknowledges financial support from PNRR MUR NQSTI PE000000023 project HIQRES CUP H43C22000870001.